1. Determine the area under the normal curve between Z = -1.13 and Z = .67

2. Determine the area under the normal curve for Z < -.23, plus Z > 1.38 (sum of each area)

Please someone help me with these two problem and explain how to do it ?

Thank you !!!
Answer #1

Using standard normal table ,

1)

P(-1.13 < z < 0.67)

= P(z < 0.67) - P(z < -1.13)

= 0.6193

Area = 0.6193

2)

P(z < -0.23) = 0.4090

P(Z > 1.38) = 0.352

Area = 0.4090 + 0.352 = 0.7610